Please Help soon as possible!

logx+log3 = log30
Thank you.

Answer by josmiceli(6783)

To find the log of a product, add the logs
log(a x b) = log(a) + log(b)
log(x) + log(3) = log(30)
log(30x) = log(30)
30x = 30
x = 10